OpenStudy (anonymous):

x^2+y^2=4 1)express each relation as an explicit function of x. 2) derivative

OpenStudy (anonymous):

i.e. "solve for \(y\)" \[x^2+y^2=4\] a circle with center \((0,0)\) and radius \(2\) so not a function, but you can solve via \[y^2=4-x^2\] \[y=\pm\sqrt{4-x^2}\]

OpenStudy (anonymous):

I got that but I don't know how to do the second part

myininaya (myininaya):

You will have two functions so you will have two derivatives, one for y>=0 and the other for y<=0.

myininaya (myininaya):

\[y=\sqrt{4-x^2} ; y=-\sqrt{4-x^2} \] Find derivative of both of these by chain rule.

OpenStudy (jonnyvonny):